Derive the expression or time period of a satelite revolving about its plane

Grade:11

1 Answers

Monika
6 years ago
Time period of a satellite Time taken by the satellite to complete one revolution round the Earth is called time period.Time period, T = circumference of the orbit / orbital velocityT = 2πr / v0 = 2π(R +h) / v0where r is the radius of the orbit which is equal to (R+h).v0 = root[GM/R+h] soT = 2π (R+h) root[(G+h)/GM]As GM = gR2, T = 2π root[(R+h)3/gR2]If the satellite orbits very close to the Earth, then h